# CoinFT: A Coin-Sized, Capacitive 6-Axis Force Torque Sensor for Robotic Applications

### v2

@@ Abstract
We introduce CoinFT, a capacitive 6-axis force / torque (F / T) sensor that is compact, light, low-cost, and robust with an average mean-squared error of 0.11 N for force and 0.84 mNm for moment when the input ranges from 0$\sim$10 N and 0$\sim$4 N in normal and shear directions, respectively. CoinFT is a stack of two rigid PCBs with comb-shaped electrodes connected by an array of silicone rubber pillars. The microcontroller interrogates the electrodes in different subsets in order to enhance sensitivity for measuring 6-axis F / T. The combination of desirable features of CoinFT enables various contact-rich robot interactions at a scale, across different embodiment domains including drones, robot end-effectors, and wearable haptic devices. We demonstrate the utility of CoinFT on drones by performing an attitude-based force control to perform tasks that require careful contact force modulation. The design, fabrication, and firmware of CoinFT are open-sourced at https://hojung-choi.github.io/coinft.github.io/ .
@@ I Introduction
In this paper, we introduce CoinFT, a capacitive 6-axis F / T sensor the size of a US quarter dollar, weighing 2 g, with a material cost of less than \$ 10 (Fig. 1 (a)). CoinFT combines the aforementioned desirable features, unlocking various contact-rich robot interactions in unstructured environments.
One particularly promising application for CoinFT is on drones performing forceful interactions with the environment. Drones have previously executed contact-rich tasks without force feedback [11] in controlled scenarios or through careful human teleoperation [6]. While force sensors have been demonstrated in such tasks, their size, weight, and fragility make them unsuitable for unstructured, outdoor environments [12]. CoinFT’s design addresses these challenges, making it a viable solution for drones operating in environments where large impulses from contact or crashes are common. Its lightweight nature minimally impacts drone operation time, and its compact size facilitates easy integration into even small drones. In this work, we demonstrate the use of CoinFT on a drone for contact modulation through PID force control and the deployment of objects onto environmental surfaces by applying a controlled range of force.
@@
  • Experimental validation of the feasibility of integrating CoinFT on drones by modulating contact force and deploying objects on environment surfaces.
@@ II-A Multi-axial Force Torque Sensors
Capacitive sensing is another widely adopted approach for designing 6-axis F / T sensors, valued for its potential to achieve a compact form factor while maintaining high sensitivity [8]. Compared to piezoresistive sensors, capacitive sensors can be relatively simpler to fabricate. Multiaxial F / T sensing via capacitive methods is typically achieved in one of two ways: using a single layer of sensing electrodes integrated with specially designed structures that deform and create distinct signal patterns upon input [15, 16, 17], or employing multi-layer structures with electrodes, where each layer is sensitive to specific subsets of forces and torques, generating unique signal patterns accordingly [18].
@@
Optical sensing methods have been employed in the design of multi-axis F / T sensors by controlling the scattering or reflection of light within the system [19]. These methods can be broadly categorized into light intensity sensing [20], employing Fiber Bragg Grating (FBG) optical fibers [21], and vision-based techniques. In traditional optical methods, variations in reflected light intensity at multiple locations are converted into measurements of force and torque [20]. FBG-based sensing leverages FBG fibers embedded in areas of higher strain within the sensor body, functioning similarly to strain gauges to precisely measure deformation [21]. Vision-based tactile sensing involves using a compact camera to observe the deformation of a silicone rubber medium, often a robot fingertip, upon contact [22, 23]. A recently popularized approach for force and torque sensing involves monitoring the deformation of specially patterned silicone rubber fingertips and employing data-driven models to infer the applied force and torque [23, 24, 25]. Another method involves tracking the movement of a visual cue with a camera to deduce the load [26, 27]. Alternatively, gel deformation patterns can be remotely observed through an optical fiber, enabling the decoupling of the silicone gel from the sensing camera [28].
@@ III-A CoinFT Design
CoinFT is a capacitive 6-axis force / torque sensor. Its circular sensing area is 20 mm in diameter and approximately 2 mm in thickness (Fig. 1 (c)), having a comparable form-factor to a quarter-dollar coin in the United States currency (Fig. 1 (a)). With a total weight of 2 g, it consists of two 0.8 mm thick rigid PCBs (Printed Circuit Boards) connected by an array of 127 um tall cylindrical pillars cast from silicone rubber (TAP® Silicone RTV Mold-Making System). The radial density of pillars increases towards the edge of the PCB for improved strength against delamination. We employ a dense array of microns-wide pillars, despite the added challenges in fabrication, to retain sensitivity. Wider structures will increase effective stiffness of the pillar layer, which will be discussed in detail in Section III-D.
Figure 1: (a) CoinFT is approximately the size of a U.S. quarter-dollar coin. (b) Exploded view of CoinFT. It consists of two rigid PCBs (upper and lower sensing layers) connected with an array of silicone rubber pillars. The fPCB top shield layer provides passive shielding. (c) Dimensions of each layer. The overall thickness is approximately 2 mm.
Figure 2: CoinFT working principle. The microcontroller firmware is programmed to switch the pair of rigid PCBs between two different configurations - normal mode and shear mode. The capacitance signals collected from these two configurations produce unique patterns under different force and torque inputs.
Each rigid PCB is multilayered, with four quadrants of a pair of comb-shaped electrodes on one side and a plane electrode for passive shielding on the other side (Fig. 1 (b), (c)). The upper sensing layer has eight physical electrodes while the lower sensing layer houses two physical electrodes (Fig. 1 (b)). It is by leveraging this design with the unique capabilities of the microcontroller (Cypress, PSoCTM 4100S) on CoinFT, that a pair of PCBs becomes sensitive to six axes of force and torque, which will be discussed in detail in Section III-B. The top shield layer made of a 127 um thick fPCB (flexible Printed Circuit Board) is then added to provide additional passive shielding, as the via holes in the upper sensing layer can introduce noise from stray capacitance in the environment (Fig. 1 (b), (c)). Each PCB layer has structures for alignment on its sides, which are later removed during the final stages of fabrication.
A tab (10 mm $\times$ 11 mm) attached to the side of the circular sensing area houses the electronics necessary for capacitance sensing (Fig. 1 (b), (c)). A microcontroller with external capacitors measures and converts capacitance to digital counts using its native capacitance-to-digital conversion (CDC) module at 360 Hz. Signals can be acquired by other devices through direct serial communication or USB ports through a serial to USB converter, such as a PSoCTM KitProg2, with a latency of 2 ms on average.
@@ III-B Working Principle
In the shear mode, electrodes in the upper sensing layer are kept as independent sensing electrodes ($X1$, $X2$, $X3$, $X4$, $Y1$, $Y2$, $Y3$, and $Y4$), while the electrodes in the lower sensing layer are configured such that the ground and shield electrodes form an alternating pattern (Fig. 2). The shield electrode actively mitigates interference from stray capacitance by being excited to the same electrical potential as the sensing electrodes, an innate capability of PSoCTM 4100S. This design allows the shear mode to be sensitive primarily to shear inputs such as $Fx$, $Fy$, and $Tz$. Under a shear pressure, the relative position of the upper sensing layer and the lower layer shifts laterally, changing $A$ between a sensing electrode and ground electrode. Due to the negligible capacitance formation between a sensing electrode and a shield electrode, the relative position change results in a differential signal pattern in the shear mode. The comb-shaped electrodes are populated in a way that allows the first and third quadrant of the shear mode to be sensitive to loads in the $X$ direction, while the second and fourth quadrants are sensitive to the $Y$ direction. Under torsion ($Tz$), all four quadrants in the shear mode show a differential signal pattern. Unlike the normal mode configuration where there is negligible cross-modal sensitivity by design, there is cross-coupling in the shear mode configuration. This effect and its implications will be discussed in detail Section III-D1. The unique signal pattern generated from the normal and shear mode configurations upon loads in different axes makes CoinFT a 6-axis F / T sensor.
@@ III-C Sensor Fabrication
On the upper sensing layer PCB, after the microcontroller and the electrical components are hand-soldered, the sensing area is primed and applied with vacuum degassed uncured silicone rubber (Fig. 3 (f)). Then, the silicone layer is spin-coated at 5000 rpm for 75 seconds to achieve a uniform and thin layer (Fig. 3 (g)). The pillar layer is attached to the upper sensing layer PCB with a 203 um polyimide film spacer in between, to control the distance between the two PCBs. A 100 g mass and an acrylic plate is placed on top of the assembly to apply equally distributed pressure, ensuring that the pillars and the spun coat silicone layer are in contact (Fig. 3 (h)). The assembly is then cured inside a 65 °C oven for 72 hours. After the alignment pins and spacers are removed, the top shield layer fPCB is adhered to the other side of the upper sensing layer PCB using an adhesive (Henkel Corporation, Loctite 401 Instant Adhesive) (Fig. 3 (i)). The cross-section of the pillar assembly under a microscope is shown in Fig. 3 (j). The total cost of all components is less than \$10.
@@ III-D1 CoinFT Calibration
Based on the working principle from Fig. 2, CoinFT produces a distinguishable signal pattern from its 12 sensing electrodes (4 normal mode, 8 shear mode) under different force and torque loads. The simulated sensor response (Fig. 4 (a)-(c)) shows the distinguishable signal patterns under ideal alignment and distance of the upper and lower sensing layer PCBs. For sensor calibration, CoinFT is mounted on a commercial sensor (ATI Industrial Automation, Gamma) which provides ground truth force and torque information, while a combination of different loads are given from a human finger. A 2nd order least squares fit is used to map the raw capacitance signals into corresponding force and torque. The calibration results are shown in Fig. 5, and the mean square error (MSE) is provided in Table I. The minimum detectable force of CoinFT is 20 mN.
TABLE I: Mean-squared error of CoinFT calibration results.
Input Range: 0$\sim$5 N Normal, 0$\sim$2 N Shear
@@
Normal + Shear 0.04 0.04 0.05 0.32 0.31 0.24
Shear 0.04 0.04 0.11 0.34 0.32 0.24
Input Range: 0$\sim$10 N Normal, 0$\sim$4 N Shear
@@
Normal + Shear 0.12 0.10 0.10 0.84 1.03 0.65
Shear 0.13 0.10 0.22 0.91 1.2 0.72
  • *”Normal + Shear” includes all 12 electrodes in the calibration process, while the ”Shear” case uses only 8 electrodes from the shear mode. Forces are measured in N, and moments are in mNm. A CoinFT sample with a pillar diameter of 100 um was used.
The 8 shear mode electrodes respond to inputs causing change in distance, $Fz$, $Mx$, and $My$, between the electrodes and the ground electrode (Fig. 4 (a)). While it is possible to calibrate the sensor using only the shear mode elecrodes, the redundancy added by the high signal-to-noise ratio (SNR) by the normal mode signals (Fig. 4 (a)) significantly increases calibration accuracy, as shown in Table I. As expected, the increase in accuracy is more significant for $Fz$, $Mx$, and $My$.

### v3

@@ Abstract
[…]
@@ I Introduction
In this paper, we introduce CoinFT, a capacitive 6-axis F / T sensor the size of a US quarter dollar, weighing 2 g (Fig. 1 (a)), with a material cost of less than \$ 11 (Table II). CoinFT combines the aforementioned desirable features, unlocking various contact-rich robot interactions in unstructured environments.
We validate CoinFT’s multi-axis sensing capability through a contact-probing experiment, where a 7-DoF robotic arm applies controlled 6-axis F / T on a fingertip-mounted CoinFT. This setup demonstrates CoinFT’s suitability for fine-grained manipulation scenarios while maintaining a compact fingertip-scale form factor.
Another promising application for CoinFT is on drones performing forceful interactions with the environment. Drones have previously executed contact-rich tasks without force feedback [11] in controlled scenarios or through careful human teleoperation [6]. While force sensors have been demonstrated in such tasks, their size, weight, and fragility make them unsuitable for unstructured, environments [12]. CoinFT’s design addresses these challenges, making it a viable solution for drones operating in environments where large impulses from contact or crashes are common. Its lightweight nature minimally impacts drone operation time, and its compact size facilitates easy integration into even small drones. In this work, we demonstrate the use of CoinFT on a drone for contact modulation through PID force control and the deployment of objects onto environmental surfaces by applying a controlled range of force.
@@
  • Experimental validation of CoinFT on multiple robotic platforms, including a contact-probing experiment demonstrating accurate multi-axis sensing at the fingertip scale and a drone performing contact-force modulation.
@@ II-A Multi-axial Force Torque Sensors
TABLE I: Comparison of representative lower-cost 6-axis force/torque sensors for robotic applications
Sensor Transduction Size [mm] Mass Sensing range Accuracy (R2 or % error) Bandwidth Cost Robustness
MinebeaMitsumi™ 6-axis F/T sensor Piezoresistive $\phi$ 9.6 $\times$ $t$ 9 3 g F: 40 N T: 400 mN·m Full-scale error: $\pm$5% 1 kHz \$1200+ No
Guggenheim et al. [10] Barometric (MEMS) $75\times 75\times 30$ N/A $F_{x,y}$: 5 N $F_{z}$: 12 N $T$: 3000 mN·m $F_{x}$: 0.937 $F_{y}$: 0.909 $F_{z}$: 0.998 $M_{x}$: 0.984 $M_{y}$: 0.984 $M_{z}$: 0.991 50 Hz \$20 Yes
Ouyang et al. [15] Vision (Camera) $36\times 31\times 51$ N/A $F_{x,y}$: 2 N $F_{z}$: 1 N $T_{x,y}$: 100 mN·m $T_{z}$: 4 mN·m $F_{x}$: 0.991 $F_{y}$: 0.996 $F_{z}$: 0.875 $M_{x}$: 0.997 $M_{y}$: 0.998 $M_{z}$: 0.902 25 Hz (sampling) \$40 Yes
Kim et al. [16] Capacitive $14\times 17\times 6.5$ 2.6 g F: 30 N M: 300 mN·m Full-scale error: Fx: 0.92% Fy: 0.83% Fz: 1.57% Mx: 2.8% My: 2.1% Mz: 3.4% 200 Hz (sampling) <<5 Hz (mech. estimated) N/A N/A
El-Azizi et al. [17] Optical $\phi$ 27 $\times$ $t$ 20 16.4 g F: 10 N M: 200 mN·m $F_{x}$: 0.990 $F_{y}$: 0.993 $F_{z}$: 0.560 $M_{x}$: 0.995 $M_{y}$: 0.991 $M_{z}$: 0.923 2.5 kHz (sampling) \$100 Yes
CoinFT (this work) Capacitive $\phi$ 20 $\times$ $t$ 2 2 g $F_{x,y}$: 5 N $F_{z}$: 15 N $T_{x,y}$: 100 mN·m $T_{z}$: 80 mN·m $F_{x}$: 0.983 $F_{y}$: 0.983 $F_{z}$: 0.998 $M_{x}$: 0.988 $M_{y}$: 0.987 $M_{z}$: 0.989 360 Hz (sampling) 100 Hz (mech.) \$11 Yes
@@
Capacitive sensing is another widely adopted approach for designing 6-axis F / T sensors, valued for its potential to achieve a compact form factor while maintaining high sensitivity [8]. Compared to piezoresistive sensors, capacitive sensors can be relatively simpler to fabricate. Multiaxial F / T sensing via capacitive methods is typically achieved in one of two ways. The first uses a single layer of sensing electrodes integrated with specially designed structures that deform to produce distinct signal patterns upon loading [18, 16, 19]. However, such single-layer designs often require complex fabrication processes and specialized equipment for custom machining or microstructuring, which limits scalability. The second approach employs multi-layer electrode stacks, where each layer responds to specific subsets of forces and torques to generate distinguishable signal patterns [20]. While effective, these architectures introduce additional fabrication complexity and can be bulkier, making them less suitable for applications that require compactness. The CoinFT design draws inspiration from multi-layer capacitive architectures but achieves comparable multi-axis capability using a single pair of sensing electrodes through an efficient mode-switching and sampling strategy, as discussed in Section III-A.
@@
Optical sensing methods have been employed in the design of multi-axis F / T sensors by controlling the scattering or reflection of light within the system [21]. These methods can be broadly categorized into light intensity sensing [22], employing Fiber Bragg Grating (FBG) optical fibers [23], and vision-based techniques. In traditional optical methods, variations in reflected light intensity at multiple locations are converted into measurements of force and torque [22]. FBG-based sensing leverages FBG fibers embedded in areas of higher strain within the sensor body, functioning similarly to strain gauges to precisely measure deformation [23]. Vision-based tactile sensing involves using a compact camera to observe the deformation of a silicone rubber medium, often a robot fingertip, upon contact [24, 25, 26]. A recently popularized approach for force and torque sensing involves monitoring the deformation of specially patterned silicone rubber fingertips and employing data-driven models to infer the applied force and torque [25, 27, 28]. Another method involves tracking the movement of a visual cue with a camera to deduce the load [29, 30]. Alternatively, gel deformation patterns can be remotely observed through an optical fiber, enabling the decoupling of the silicone gel from the sensing camera [31].
@@
Among the various 6-axis F/T sensors reported in prior work, Table I compares representative designs that emphasize compact form factors and relatively low cost, including both research prototypes and commercial devices. CoinFT combines several desirable attributes such as compact and low-profile design, reasonable accuracy across all six axes, mechanical robustness, and affordability, within a single platform. This balance of performance and practicality enables a wide range of robotic applications, including wearable haptic interfaces[7, 35], dexterous manipulation[36], and drone-based interaction tasks, as demonstrated in Section IV-B.
@@ III-A CoinFT Design
[…]
Figure 1: (a) CoinFT is approximately the size of a U.S. quarter-dollar coin. (b) Exploded view of CoinFT. It consists of two rigid PCBs (upper and lower sensing layers) connected with an array of silicone rubber pillars. The fPCB top shield layer provides passive shielding. The alignment tabs are removed after assembly. (c) Dimensions of each layer. The overall thickness is approximately 2 mm.
Figure 2: CoinFT working principle. The microcontroller firmware is programmed to switch the pair of rigid PCBs between two different electrode configurations - normal mode and shear mode. The capacitance signals collected from these two configurations produce unique patterns under different force and torque inputs.
Each rigid PCB is multilayered, with four quadrants of a pair of comb-shaped electrodes on one side and a plane electrode for passive shielding on the other side (Fig. 1 (b), (c)). The upper sensing layer has eight physical electrodes while the lower sensing layer houses two physical electrodes (Fig. 1 (b)). It is by leveraging this design with the unique capabilities of the microcontroller (Cypress, PSoCTM 4100S) on CoinFT, that a pair of PCBs becomes sensitive to six axes of force and torque, which will be discussed in detail in Section III-B. The top shield layer made of a 127 um thick fPCB (flexible Printed Circuit Board) is then added to provide additional passive shielding, as the via holes in the upper sensing layer can introduce noise from stray capacitance in the environment (Fig. 1 (b), (c)). Each PCB layer has an alignment tab on its sides, which are later removed during the final stages of fabrication.
An electronics tab (10 mm $\times$ 11 mm) attached to the side of the circular sensing area houses the electronics necessary for capacitance sensing (Fig. 1 (b), (c)). A microcontroller with external capacitors measures and converts capacitance to digital counts using its native capacitance-to-digital conversion (CDC) module at 360 Hz. Signals can be acquired by other devices through direct serial communication or USB ports through a serial to USB converter, such as a PSoCTM KitProg2, with a latency of 2 ms on average.
@@ III-B Working Principle
In the shear mode, electrodes in the upper sensing layer are kept as independent sensing electrodes ($X1$, $X2$, $X3$, $X4$, $Y1$, $Y2$, $Y3$, and $Y4$), while the electrodes in the lower sensing layer are configured such that the ground and shield electrodes form an alternating pattern (Fig. 2). The shield electrode actively mitigates interference from stray capacitance by being excited to the same electrical potential as the sensing electrodes, an innate capability of PSoCTM 4100S. This design allows the shear mode to be sensitive primarily to shear inputs such as $Fx$, FyFy, and $Tz$. Under a shear pressure, the relative position of the upper sensing layer and the lower layer shifts laterally, changing $A$ between a sensing electrode and ground electrode. Due to the negligible capacitance formation between a sensing electrode and a shield electrode, the relative position change results in a differential signal pattern in the shear mode. The comb-shaped electrodes are populated in a way that allows the first and third quadrant of the shear mode to be sensitive to loads in the $X$ direction, while the second and fourth quadrants are sensitive to the $Y$ direction. Under torsion ($Tz$), all four quadrants in the shear mode show a differential signal pattern. Unlike the normal mode configuration where there is negligible cross-modal sensitivity by design, there is cross-coupling in the shear mode configuration. This is resolved in our calibration, as discussed in detail in Section III-D1. The unique signal pattern generated from the normal and shear mode configurations upon loads in different axes makes CoinFT a 6-axis F / T sensor.
@@ III-C Sensor Fabrication
TABLE II: Bill of Materials (BOM) for CoinFT
Item Quantity Unit Cost [\$] Total Cost [\$]
PSoC 4100S 1 3.25 3.25
Middle layer PCB 1 2.94 2.94
Top layer PCB 1 0.76 0.76
Shield PCB 1 2.09 2.09
SMD electronics 6 0.20 1.20
Molex connector 1 0.69 0.69
Total 10.93
On the upper sensing layer PCB, after the microcontroller and the electrical components are hand-soldered, the sensing area is primed and applied with vacuum degassed uncured silicone rubber (Fig. 3 (f)). Then, the silicone layer is spin-coated at 5000 rpm for 75 seconds to achieve a uniform and thin layer (Fig. 3 (g)). The pillar layer is attached to the upper sensing layer PCB with a 203 um polyimide film spacer in between, to control the distance between the two PCBs. A 100 g mass and an acrylic plate is placed on top of the assembly to apply equally distributed pressure, ensuring that the pillars and the spun coat silicone layer are in contact (Fig. 3 (h)). The assembly is then cured inside a 65 °C oven for 72 hours. After the alignment pins and spacers are removed, the top shield layer fPCB is adhered to the other side of the upper sensing layer PCB using an adhesive (Henkel Corporation, Loctite 401 Instant Adhesive) (Fig. 3 (i)). The cross-section of the pillar assembly under a microscope is shown in Fig. 3 (j). The total cost of all components is approximately \$11, without scaled production (Table II).
@@ III-D1 CoinFT Calibration
Based on the working principle from Fig. 2, CoinFT produces a distinguishable signal pattern from its 12 sensing electrodes (4 normal mode, 8 shear mode) under different force and torque loads. The simulated sensor response (Fig. 4 (a)-(c)) shows the distinguishable signal patterns under ideal alignment and distance of the upper and lower sensing layer PCBs. For sensor calibration, CoinFT is mounted on a commercial sensor (ATI Industrial Automation, Gamma) which provides ground truth force and torque information, while a combination of different loads are given from a human finger. The calibration procedure is purely data-driven. Ten trials, each consisting of 35 seconds of data, are used for training, and one additional trial of the same length is reserved for testing. A second-order least-squares regression is employed to map the raw capacitance signals to the corresponding force and torque values. Let $X\in\mathbb{R}^{24\times N}$ denote the feature matrix containing the first- and second-order capacitance terms from $N$ samples, and let $Y\in\mathbb{R}^{6\times N}$ denote the corresponding reference forces and torques. The calibration matrix $A\in\mathbb{R}^{6\times 24}$ is obtained by solving the normal equation. The calibration results on the test set is shown in Fig. 5, and the root mean square error (RMSE) and $R^{2}$ values are provided in Table III. The minimum detectable force of CoinFT is 20 mN.
TABLE III: Accuracy of CoinFT calibration.
[…]
@@
Normal+Shear (RMSE) 0.038 0.040 0.050 0.317 0.306 0.231
Shear (RMSE) 0.039 0.040 0.115 0.343 0.323 0.244
Normal+Shear ($R^{2}$) 0.994 0.994 0.997 0.995 0.995 0.992
Shear ($R^{2}$) 0.993 0.994 0.986 0.994 0.994 0.992
[…]
@@
Normal+Shear (RMSE) 0.186 0.134 0.186 1.114 1.322 0.810
Shear (RMSE) 0.174 0.129 0.266 1.169 1.587 0.867
Normal+Shear ($R^{2}$) 0.983 0.983 0.998 0.988 0.987 0.989
Shear ($R^{2}$) 0.979 0.985 0.996 0.986 0.981 0.986
  • *”Normal + Shear” includes all 12 electrodes in the calibration process, while the ”Shear” case uses only 8 electrodes from the shear mode. Forces errors are reported in $N$, and moment errors are reported in $mNm$. A CoinFT sample with a pillar diameter of 100 um was used. RMSE stands for Root Mean Squared Error.
The 8 shear mode electrodes respond to inputs causing change in distance, $Fz$, MxMx, and $My$, between the electrodes and the ground electrode (Fig. 4 (a)). While it is possible to calibrate the sensor using only the shear mode electrodes, the redundancy added by the high signal-to-noise ratio (SNR) by the normal mode signals (Fig. 4 (a)) significantly increases calibration accuracy, as shown in Table III. As expected, the increase in accuracy is more significant for $Fz$, MxMx, and $My$.